# Fixed-Point RNNs: Interpolating from Diagonal to Dense

### v2

@@ Abstract
Linear recurrent neural networks (RNNs) and state-space models (SSMs) such as Mamba have become promising alternatives to softmax-attention as sequence mixing layers in Transformer architectures. Current models, however, do not exhibit the full state-tracking expressivity of RNNs because they rely on channel-wise (i.e. diagonal) sequence mixing. In this paper, we investigate parameterizations of a large class of dense linear RNNs as fixed-points of parallelizable diagonal linear RNNs. The resulting models can naturally trade expressivity for efficiency at a fixed number of parameters and achieve state-of-the-art results on the commonly used toy tasks $A_{5}$, $S_{5}$, copying, and modular arithmetics.
Figure 1: Sequence length generalization at training length 16 (pink) for state-tracking on $A_{5}$, with Transformer (brown) and LSTM (purple) as lower/upper bounds. Our Fixed-Point RNN (FP-Mamba) is trained at different maximum number of fixed-point iterations $\ell_{\text{max}}$: between $2$ (green) and $16$ (blue). Increasing the number of fixed-point iterations allows the linear RNN to interpolate from diagonal to dense in a few iterations.
@@ 1 Introduction
State-space models (SSMs) and other new efficient recurrent token mixers are becoming a popular alternative to softmax attention in language modeling (Gu & Dao, 2023) as well as in other applications such as vision (Liu et al., 2024) and DNA processing (Nguyen et al., 2024). Inspired by linear input-controlled filtering, these models can be expressed as carefully parametrized linear recurrent neural networks (RNNs) with input-dependent, diagonal state transition:
$h_{t}=\text{diag}(\mathbf{a}_{t})h_{t-1}+\mathbf{B}_{t}x_{t}$ (1)
Compared to classical RNNs such as LSTMs (Hochreiter & Schmidhuber, 1997), in Eq. (1) the relation between the previous hidden state $h_{t-1}$ and the current $h_{t}$ is linear and its coefficient $\mathbf{a}_{t}$ does not depend on the hidden states. These choices allow SSMs such as Mamba (Gu & Dao, 2023) to be computed through efficient parallel methods during training. Furthermore, they are easier to optimize than classical RNNs, thanks to stable and efficient reparametrizations available for diagonal transitions (Orvieto et al., 2023; Zucchet & Orvieto, 2024) – techniques that are significantly more difficult to apply effectively in the classical setting (Arjovsky et al., 2016; Helfrich et al., 2018). At test time, they are faster than classical Transformers on long sequences due to their recurrent nature.
Though modern linear RNNs have shown promise in practice, recent theoretical studies suggest that using dense, input-dependent transition matrices (i.e. replacing $\text{diag}(\mathbf{a}_{t})$ with a dense $\mathbf{A}_{t}$) could present an opportunity to improve expressivity and unlock performance on challenging tasks. In particular, Cirone et al. (2024b) prove that dense selective SSMs are endowed with the theoretical expressivity of classical non-linear RNNs such as LSTMs. As shown by Merrill et al. (2024) and Sarrof et al. (2024), such gained expressivity proves to be particularly useful in state-tracking applications where models are expected to maintain and extrapolate a complex state of the world. Since state-tracking is naturally expressed by non-linear RNNs but provably unavailable to channel-wise sequence mixers such as SSMs or Transformers, Merrill & Sabharwal (2023a) speculate on a fundamental tradeoff between parallelism and expressivity. This discussion sparked interest in non-diagonal recurrences and parallelizable architectures capable of state-tracking (Grazzi et al., 2024; Terzic et al., 2025; Schöne et al., 2025; Peng et al., 2025; Siems et al., 2025).
When designing new architectures involving dense selective yet linear state transitions of the form $h_{t}=\mathbf{A}_{t}h_{t-1}+\mathbf{B}_{t}x_{t}$, two fundamental concerns arise:
@@
    What should the parametric form for $\mathbf{A}_{t}$, as a function of the input be? How can we guarantee this parametrization induces a stable recurrence, like in standard [^1] Standard SSMs are diagonal and operate in polar coordinates, parametrizing directly the gap between eigenvalues and the stability threshold (Orvieto et al., 2023). This technique allows to increasing granularity near the identity, and to effectively normalize the forward pass (cf. $\sqrt{1-|\gamma|^{2}}$ term in Griffin (De et al., 2024)). SSMs?
@@
Perhaps the first approach tackling the above questions was DeltaNet (Schlag et al., 2021; Yang et al., 2024b) with a block-diagonal and orthogonal therefore, stable state transition structure, where each block is parametrized by a Householder matrix. The parallelizable algorithm, was then extended to include negative eigenvalues (Grazzi et al., 2024), gates (Yang et al., 2025), and most recently products of Householders (Siems et al., 2025). Such choices, leading to increased expressivity as exemplified by their state-tracking and length generalization capabilities, are motivated mainly by hardware considerations: Householder-based mixing can be implemented efficiently on GPUs as linear attention via WY-representations and the UT transform (Yang et al., 2024b).
Figure 2: (a) State-tracking on $A_{5}$ at sequence length $16$, and (b) character accuracy of copying at 2×2$\times$2 × sequence length generalization, trained on lengths $\in[5,50]$. Our single layer FP-Mamba with mixer reflections $r\in\{1,2,4\}$ is compared to baselines of increasing depth $\in\{1,2,4,6,8\}$. FP-Mamba is the only model capable of solving both the state-tracking and the copy task.
@@
Motivated by the idea of designing a parallelizable general-purpose method to implement new dense RNN variations, in this paper we devise a new adaptive computation strategy which allows to interpolate between fast recurrent diagonal RNNs and dense recurrences with arbitrary preselected structure. Instead of parametrizing the dense RNN layer as an explicit function $h=F_{\theta}(x)$, we build on the literature of equilibrium/implicit models (Bai et al., 2019; Ghaoui et al., 2021) to parametrize it implicitly as a solution $h^{*}$ to a fixed-point equation $h=f_{\theta}(x,h)$ involving only a diagonal RNN. As described in Fig. 9(a), we solve for $h^{*}$ using a fixed-point iteration of diagonal RNN evaluations $f_{\theta}$.
@@ 2 Background
Since their introduction (Rumelhart et al., 1986; Elman, 1990), RNNs have significantly contributed to the evolution of machine learning methods for sequential data (Hochreiter & Schmidhuber, 1997; Jaeger, 2001). But despite their theoretical promise of Turing-completeness (Siegelmann & Sontag, 1992), recurrent models fell out of fashion due to two significant challenges: they are inherently sequential, and notoriously difficult to train (Hochreiter et al., 2001; Pascanu et al., 2013). The recent advancements of linear RNNs (Gu & Dao, 2023) suggest a way forward to combine the scalability of Transformers (Vaswani et al., 2017) with the expressivity of classical RNNs (Cirone et al., 2024b). The key challenge here is the stable and efficient parametrization of a linear RNN layer with a time-varying recurrent transition matrix. In this paper we are exploring first steps towards this goal.
@@
$F_{\theta}:x\mapsto h,$ $h_{t}=\mathbf{A}_{t}h_{t-1}+\mathbf{B}_{t}x_{t},$ (2)
where $\mathbf{A}_{t}\in\mathbb{R}^{d\times d}$ corresponds to the time-varying state transition matrix, $\mathbf{B}_{t}\in\mathbb{R}^{d\times d}$ is the input transformation matrix, $h_{t}\in\mathbb{R}^{d}$ denotes the hidden state, and $x_{t}\in\mathbb{R}^{d}$ is the input for $t<T$ steps. For a given sequence of $\mathbf{A}_{t}$, the complexity of a forward pass is $O(Td^{2})$ in memory and $O(T)$ sequential steps. Although such a linear RNN could also be computed in $O(\log T)$ sequential steps using a parallel scan algorithm (Martin & Cundy, 2018), this would require materializing matrix-matrix multiplications at cost $O(d^{3})$. An issue in both scenarios, however, is the parametrization of $\mathbf{A}_{t}$ as time-varying, i.e. input- or even hidden state-dependent matrices. In general, this requires a map $\mathcal{M}:d\mapsto d^{2}$ with potentially $d\times d^{2}$ parameters, and $O(Td^{3})$ time complexity. While structured dense matrix representations for $\mathbf{A}_{t}$ could potentially present a remedy, they come with additional challenges: (1) In order to guarantee expressivity, the $\mathbf{A}_{t}$ cannot be co-diagonalizable such as for example Toeplitz matrices (Cirone et al., 2024b). (2) In order to guarantee stability of the dynamical system, the spectral radius $\rho(\mathbf{A}_{t})$ needs to be less than, but still close to $1$ for long-range interactions (Orvieto et al., 2023). (3) The matrix structure needs to be closed under multiplications to enable parallel scans without having to materialize dense representations at $O(Td^{2})$ memory cost.
@@ 3 Fixed-Points as an RNN Layer
In this section, we introduce an implicit parameterization for a family of dense RNNs $F_{\theta}(x)$ which describes its output by a solution $h^{*}\in\mathbb{R}^{T\times d}$ to the fixed-point equation $h=f_{\theta}(x,h)$ (Sec. 3.1). Then, we discuss how to find the solution $h^{*}$ using fixed-point iterations (Sec. 3.2) and the algorithmic implications (Sec. 3.4) of the FP-RNNs framework in light of the challenges outlined in Sec. 2. Finally, we briefly touch on how to train an implicitly dense model $F_{\theta}(x)$ with gradient descent (Sec. 3.5).
@@ 3.1 From Explicit to Implicit Parameterization
We start by designing a diagonal RNN $f_{\theta}(x,h)$ such that the solution $h^{*}$ to its fixed-point equation $h=f_{\theta}(x,h)$ implicitly represents a dense RNN $h^{*}=F_{\theta}(x)$. Consider the factorized parametrization of $\mathbf{A}_{t}$ similar to the one introduced by Helfrich et al. (2018) for non-linear and time-invariant RNN:
$F_{\theta}:x\mapsto h^{*},$ $h^{*}_{t}=\mathbf{Q}^{-1}_{t}\mathbf{\Lambda}_{t}h^{*}_{t-1}+\mathbf{B}_{t}x_{t}.$ (3)
Separating $\mathbf{A}_{t}$ into a diagonal matrix $\mathbf{\Lambda}_{t}\in\mathbb{R}^{d\times d}$ and a non-diagonal invertible mixing matrix $\mathbf{Q}_{t}\in\mathbb{R}^{d\times d}$ allows to describe $h^{*}$ by only a diagonal transition $\mathbf{\Lambda}_{t}$ by reformulating Eq. 3 to
$h_{t}^{*}$ $=\mathbf{\Lambda}_{t}h_{t-1}^{*}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{*}.$ (4)
This means that the states $h^{*}=F_{\theta}(x)$ of the dense linear RNN can be implicitly described by the fixed-point $h^{*}=f_{\theta}(x,h^{*})$ of a corresponding diagonal linear RNN of the following form:
$f_{\theta}:(x,h)\mapsto h^{\prime},$ $h^{\prime}_{t}=\mathbf{\Lambda}_{t}h^{\prime}_{t-1}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}.$ (5)
In other words, if we could find the fixed-point $h^{*}=f_{\theta}(x,h^{*})\in\mathbb{R}^{T\times d}$ for the diagonal RNN defined in Eq. 5, then $h^{*}$ would describe the states of a corresponding dense RNN $h^{*}=F_{\theta}(x)$. Motivated by this insight, in Sec. 3.2 we carefully parametrize the diagonal RNN $f_{\theta}(x,h)$ and its channel mixer $\mathbf{Q}_{t}$ such that a computable fixed-point exists.
@@
Figure 3: (a) An overview of the proposed Fixed-Point RNN framework in Sec. 3. A diagonal RNN $f_{\theta}$ consisting of a sequence mixer $\mathbf{\Lambda}_{t}$ and a channel mixer $\mathbf{Q}_{t}$ is iterated until convergence towards the hidden states of an implicitly dense RNN $F_{\theta}$. (b) FP-RNN variants with channel mixer introduced in Sec. 3.3 and 3.4 solve the state-tracking task $A_{5}$ up to various sequence lengths. (c) FP-RNNs adapt their computation time to the difficulty of the task by varying the number fixed-point iterations $\ell^{*}$

### v3

@@ Abstract
Linear recurrent neural networks (RNNs) and state-space models (SSMs) such as Mamba have become promising alternatives to softmax-attention as sequence mixing layers in Transformer architectures. Current models, however, do not exhibit the full state-tracking expressivity of RNNs because they rely on channel-wise (i.e. diagonal) sequence mixing. In this paper, we investigate parameterizations of a large class of dense linear RNNs as fixed-points of parallelizable diagonal linear RNNs. The resulting models can naturally trade expressivity for efficiency at a fixed number of parameters and achieve state-of-the-art results on the state-tracking benchmarks $A_{5}$ and $S_{5}$, while matching performance on copying and other tasks.
Figure 1: Sequence length generalization at training length 16 (pink) for state-tracking on $A_{5}$, with Transformer (brown) and LSTM (purple) as lower/upper bounds. Our Fixed-Point RNN (FP-Mamba-H) is trained at different maximum number of fixed-point iterations $\ell_{\text{max}}$: between $2$ (green) and $16$ (blue). Increasing the number of fixed-point iterations allows the linear RNN to interpolate from diagonal to dense in a few iterations.
@@ 1 Introduction
State-space models (SSMs) and other new efficient recurrent token mixers are becoming a popular alternative to softmax attention in language modeling (Gu & Dao, 2024) as well as in other applications such as vision (Liu et al., 2024) and DNA processing (Nguyen et al., 2024). Inspired by linear input-controlled filtering, these models can be expressed as carefully parametrized linear recurrent neural networks (RNNs) with input-dependent, diagonal state transition:
$\mathbf{h}_{t}=\text{diag}(\mathbf{a}_{t})\mathbf{h}_{t-1}+\mathbf{B}_{t}\mathbf{x}_{t}$ (1)
Compared to classical RNNs such as LSTMs (Hochreiter & Schmidhuber, 1997), in Eq. (1) the relation between the previous hidden state $\mathbf{h}_{t-1}$ and the current $\mathbf{h}_{t}$ is linear and its coefficient $\mathbf{a}_{t}$ does not depend on the hidden states. These choices allow SSMs such as Mamba (Gu & Dao, 2024) to be computed through efficient parallel methods during training. Furthermore, they are easier to optimize than classical RNNs, thanks to stable and efficient reparametrizations available for diagonal transitions (Orvieto et al., 2023; Zucchet & Orvieto, 2024) – techniques that are significantly more difficult to apply effectively in the classical setting (Arjovsky et al., 2016; Helfrich et al., 2018). At test time, they are faster than classical Transformers on long sequences due to their recurrent nature.
Though modern linear RNNs have shown promise in practice, recent theoretical studies suggest that using dense, input-dependent transition matrices (i.e. replacing $\text{diag}(\mathbf{a}_{t})$ with a dense $\mathbf{A}_{t}$) could present an opportunity to improve expressivity and unlock performance on challenging tasks. In particular, Cirone et al. (2024b) prove that dense selective SSMs are endowed with the theoretical expressivity of classical non-linear RNNs such as LSTMs. As shown by Merrill et al. (2024) and Sarrof et al. (2024), such gained expressivity proves to be particularly useful in state-tracking applications where models are expected to maintain and extrapolate a complex state of the world. Since state-tracking is naturally expressed by non-linear RNNs but provably unavailable to channel-wise sequence mixers such as SSMs or Transformers, Merrill & Sabharwal (2023) speculate on a fundamental tradeoff between parallelism and expressivity. This discussion sparked interest in non-diagonal recurrences and parallelizable architectures capable of state-tracking (Grazzi et al., 2024; Terzic et al., 2025; Schöne et al., 2025; Peng et al., 2025; Siems et al., 2025).
When designing new architectures involving dense selective yet linear state transitions of the form $\mathbf{h}_{t}=\mathbf{A}_{t}\mathbf{h}_{t-1}+\mathbf{B}_{t}\mathbf{x}_{t}$, two fundamental concerns arise:
@@
    What should the parametric form for $\mathbf{A}_{t}$, as a function of the input be? How can we guarantee this parametrization induces a stable recurrence, like in standard [^1] Standard SSMs are diagonal and operate in polar coordinates, parameterizing directly the gap between eigenvalues and the stability threshold (Orvieto et al., 2023). This technique allows to increasing granularity near the identity, and to effectively normalize the forward pass (cf. $\sqrt{1-|\gamma|^{2}}$ term in Griffin (De et al., 2024)). SSMs?
@@
Perhaps the first approach tackling the above questions was DeltaNet (Schlag et al., 2021a; Yang et al., 2024b) with a block-diagonal and orthogonal therefore, stable state transition structure, where each block is parametrized by a Householder matrix. The parallelizable algorithm, was then extended to include negative eigenvalues (Grazzi et al., 2024), gates (Yang et al., 2025), and most recently products of Householders (Siems et al., 2025). Such choices, leading to increased expressivity as exemplified by their state-tracking and length generalization capabilities, are motivated mainly by hardware considerations: Householder-based mixing can be implemented efficiently on GPUs as linear attention via WY-representations and the UT transform (Yang et al., 2024b).
Figure 2: (a) State-tracking on $A_{5}$ at sequence length $16$, and (b) character accuracy of copying at $2\times$ sequence length generalization, trained on lengths $\in[5,50]$. Our single layer FP-Mamba-H with mixer reflections $r\in\{1,2,4\}$ is compared to baselines of increasing depth $\in\{1,2,4,6,8\}$. FP-Mamba-H is the only model capable of solving both the state-tracking and the copy task.
@@
Motivated by the idea of designing a parallelizable general-purpose method to implement new dense RNN variations, in this paper we devise a new adaptive computation strategy which allows to interpolate between fast recurrent diagonal RNNs and dense recurrences with arbitrary preselected structure. Instead of parameterizing the dense RNN layer as an explicit function $\mathbf{h}=F_{\boldsymbol{\theta}}(\mathbf{x})$, we build on the literature of equilibrium/implicit models (Bai et al., 2019; Ghaoui et al., 2021) to parametrize it implicitly as a solution $\mathbf{h}^{*}$ to a fixed-point equation $\mathbf{h}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ involving only a diagonal RNN. As described in Fig. 3(a), we solve for $\mathbf{h}^{*}$ using a fixed-point iteration of diagonal RNN evaluations $f_{\boldsymbol{\theta}}$.
@@ 2 Background
Since their introduction (Rumelhart et al., 1986; Elman, 1990), RNNs have significantly contributed to the evolution of machine learning methods for sequential data (Hochreiter & Schmidhuber, 1997; Jaeger, 2001). But despite their theoretical promise of Turing-completeness (Siegelmann & Sontag, 1992), recurrent models fell out of fashion due to two significant challenges: they are inherently sequential, and notoriously difficult to train (Hochreiter et al., 2001; Pascanu et al., 2013). The recent advancements of linear RNNs (Gu & Dao, 2024) suggest a way forward to combine the scalability of Transformers (Vaswani et al., 2017) with the expressivity of classical RNNs (Cirone et al., 2024b). The key challenge here is the stable and efficient parametrization of a linear RNN layer with a time-varying recurrent transition matrix. In this paper we are exploring first steps towards this goal.
@@
$F_{\boldsymbol{\theta}}:\mathbf{x}\mapsto\mathbf{h},$ $\mathbf{h}_{t}=\mathbf{A}_{t}\mathbf{h}_{t-1}+\mathbf{B}_{t}\mathbf{x}_{t},$ (2)
where $\mathbf{A}_{t}\in\mathbb{R}^{d\times d}$ corresponds to the time-varying state transition matrix, $\mathbf{B}_{t}\in\mathbb{R}^{d\times d}$ is the input transformation matrix, $\mathbf{h}_{t}\in\mathbb{R}^{d}$ denotes the hidden state, and $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is the input for $t<T$ steps. For a given sequence of $\mathbf{A}_{t}$, the complexity of a forward pass is $O(Td^{2})$ in memory and $O(T)$ sequential steps. Although such a linear RNN could also be computed in $O(\log T)$ sequential steps using a parallel scan algorithm (Martin & Cundy, 2018), this would require materializing matrix-matrix multiplications at cost $O(d^{3})$. An issue in both scenarios, however, is the parametrization of $\mathbf{A}_{t}$ as time-varying, i.e. input- or even hidden state-dependent matrices. In general, this requires a map $\mathcal{M}:d\mapsto d^{2}$ with potentially $d\times d^{2}$ parameters, and $O(Td^{3})$ time complexity. While structured dense matrix representations for $\mathbf{A}_{t}$ could potentially present a remedy, they come with additional challenges: (1) In order to guarantee expressivity, the $\mathbf{A}_{t}$ cannot be co-diagonalizable such as for example Toeplitz matrices (Cirone et al., 2024b). (2) In order to guarantee stability of the dynamical system, the spectral radius $\rho(\mathbf{A}_{t})$ needs to be less than, but still close to $1$ for long-range interactions (Orvieto et al., 2023). (3) The matrix structure needs to be closed under multiplications to enable parallel scans without having to materialize dense representations at $O(Td^{2})$ memory cost.
@@ 3 Fixed-Points as an RNN Layer
In this section, we introduce an implicit parameterization for a family of dense RNNs $F_{\boldsymbol{\theta}}(\mathbf{x})$ which describes its output by a solution $\mathbf{h}^{*}\in\mathbb{R}^{T\times d}$ to the fixed-point equation $\mathbf{h}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ (Sec. 3.1). Then, we discuss how to find the solution $\mathbf{h}^{*}$ using fixed-point iterations (Sec. 3.2) and the algorithmic implications (Sec. 3.4) of the FP-RNN framework in light of the challenges outlined in Sec. 2. Finally, we briefly touch on how to train an implicitly dense model $F_{\boldsymbol{\theta}}(\mathbf{x})$ with gradient descent (Sec. 3.5).
@@ 3.1 From Explicit to Implicit Parameterization
We start by designing a diagonal RNN $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ such that the solution $\mathbf{h}^{*}$ to its fixed-point equation $\mathbf{h}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ implicitly represents a dense RNN $\mathbf{h}^{*}=F_{\boldsymbol{\theta}}(\mathbf{x})$. Consider the factorized parametrization of $\mathbf{A}_{t}$ similar to the one introduced by Helfrich et al. (2018) for non-linear and time-invariant RNN:
$F_{\boldsymbol{\theta}}:\mathbf{x}\mapsto\mathbf{h}^{*},$ $\mathbf{h}^{*}_{t}=\mathbf{Q}^{-1}_{t}\mathbf{\Lambda}_{t}\mathbf{h}^{*}_{t-1}+\mathbf{B}_{t}\mathbf{x}_{t}.$ (3)
Separating $\mathbf{A}_{t}$ into a diagonal matrix $\mathbf{\Lambda}_{t}\in\mathbb{R}^{d\times d}$ and a non-diagonal invertible mixing matrix $\mathbf{Q}_{t}\in\mathbb{R}^{d\times d}$ allows to describe $\mathbf{h}^{*}$ by only a diagonal transition $\mathbf{\Lambda}_{t}$ by reformulating Eq. 3 to
$\mathbf{h}_{t}^{*}$ $=\mathbf{\Lambda}_{t}\mathbf{h}_{t-1}^{*}+\mathbf{Q}_{t}\mathbf{B}_{t}\mathbf{x}_{t}+(\mathbf{I}-\mathbf{Q}_{t})\mathbf{h}_{t}^{*}.$ (4)
This means that the states $\mathbf{h}^{*}=F_{\boldsymbol{\theta}}(\mathbf{x})$ of the dense linear RNN can be implicitly described by the fixed-point $\mathbf{h}^{*}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}^{*})$ of a corresponding diagonal linear RNN of the following form:
$f_{\boldsymbol{\theta}}:(\mathbf{x},\mathbf{h})\mapsto\mathbf{h}^{\prime},$ $\mathbf{h}^{\prime}_{t}=\mathbf{\Lambda}_{t}\mathbf{h}^{\prime}_{t-1}+\mathbf{Q}_{t}\mathbf{B}_{t}\mathbf{x}_{t}+(\mathbf{I}-\mathbf{Q}_{t})\mathbf{h}_{t}.$ (5)
In other words, if we could find the fixed-point $\mathbf{h}^{*}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}^{*})\in\mathbb{R}^{T\times d}$ for the diagonal RNN defined in Eq. 5, then $\mathbf{h}^{*}$ would describe the states of a corresponding dense RNN $\mathbf{h}^{*}=F_{\boldsymbol{\theta}}(\mathbf{x})$. Motivated by this insight, in Sec. 3.2 we carefully parametrize the diagonal RNN $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ and its channel mixer $\mathbf{Q}_{t}$ such that a computable fixed-point exists.
@@
Figure 3: (a) An overview of the proposed Fixed-Point RNN framework in Sec. 3. A diagonal RNN $f_{\boldsymbol{\theta}}$ consisting of a sequence mixer $\mathbf{\Lambda}_{t}$ and a channel mixer $\mathbf{Q}_{t}$ is iterated until convergence towards the hidden states of an implicitly dense RNN $F_{\boldsymbol{\theta}}$. (b) FP-RNN variants with channel mixer introduced in Sec. 3.3 and 3.4 solve the state-tracking task $A_{5}$ up to various sequence lengths. (c) FP-RNNs adapt their computation time to the difficulty of the task by varying the number of fixed-point iterations $\ell^{*}$ .